Sroute Manufacturing intends to increase capacity by overcoming a bottleneck operation by adding new equipment. Two vendors have presented proposals. The fixed costs are $60,000 for proposal A and $70,000 for proposal B. The variable cost is $13.00 for A and $11.00 for B. The revenue generated by each unit is $2200. 

a) The break-even point in units for the proposal by Vendor A units (round your response to the nearest whole number) 

b) The break-even point in units for the proposal by Vendor B = units (round your response to the nearest whole number).
Answer #1

1. FORMULA:
BEQ = FIXED COST / PRICE PER UNIT - VARIABLE COST

FIXED COST = 60000
VARIABLE COST = 13
PRICE PER UNIT = 22

BREAK EVEN UNITS = 60000 / (22 - 13) = 6666.67 OR 6667

2. FIXED COST = 70000
VARIABLE COST = 11
PRICE PER UNIT = 22

BREAK EVEN UNITS = 70000 / (22 - 11) = 6364